# The Klein-Gordon equation on asymptotically Minkowski spacetimes: causal propagators

### v1

@@ 8.2. With bound states
In this section, we prove Theorem 8.3 and Theorem 8.4. We make appropriate adjustments to the above propagator construction to include the possibility that there are bound states of the Hamiltonian $\Delta+m^{2}+V_{\pm}$ with (negative) energy bigger than 00. Such states appear as elements in the kernel of $(\widehat{P_{V}})_{\operatorname{ff}}(\tau)$ for $\tau\in(-m,m)$.
@@
$(\widehat{P_{V}})_{\operatorname{ff},+}(\tau)=\mbox{ indicial operator of $P_{V}$ at $\mathrm{NP}$},$
while $(\widehat{P_{V}})_{\operatorname{ff},-}(\tau)$ is the indicial operator at $\mathrm{SP}$. Similarly we write
@@
$(\widehat{P_{V}})_{\operatorname{ff},\pm}(\tau)w=0\implies w\in\mathcal{S}(\mathbb{R}^{n}),$
and thus by self-adjointness of $(\widehat{P_{V}})_{\operatorname{ff},\pm}(\tau)$,
(8.14) $\tau\in\operatorname{Ell}_{\operatorname{ff},\pm}(P_{V})\iff\ker((\widehat{P_{V}})_{\operatorname{ff},\pm}(\tau))=\{0\}.$
@@
It is possible that there are finitely many points $\tau_{0}\in(-m,m)$ such that $(\widehat{P_{V}})_{\operatorname{ff}}(\tau_{0})$ is not invertible, and the remainder of this section proves estimates near such $\tau_{0}\in W^{\perp}$.
@@
$E_{\pm}(\lambda(\tau_{0}))\coloneqq\ker\left((\widehat{P_{V}})_{\operatorname{ff},\pm}(\tau_{0})\colon H_{\mathrm{sc}}^{s,\ell}(\operatorname{ff})\longrightarrow H_{\mathrm{sc}}^{s-2,\ell}(\operatorname{ff})\right).$
@@
Assume that (8.1) and (8.17) hold. Then for any $\epsilon>0$, there exists $Q\in{}^{3\mathrm{sc}}\Psi^{0,0}(\mathbb{R}^{n+1})$ such that
@@
Let $\tau_{0}\in\mathcal{B}_{V,+}$, i.e. let $\tau_{0}\in(-m,m)$, $\tau_{0}\neq 0$ and $E_{+}(\lambda(\tau_{0}))\neq\{0\}$. For any $Q\in{}^{3\mathrm{sc}}\Psi^{0,0}(\mathbb{R}^{n+1})$ with $\tau_{0}\subset\operatorname{Ell}_{\operatorname{ff}}(Q)$ and ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)$ sufficiently small, for any, $G\in{}^{3\mathrm{sc}}\Psi$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$, then for $s,\ell,N,M\in\mathbb{R}$, the estimates in Proposition 8.7 hold.
@@
(8.23) $K_{\tau_{0}}\in{}^{3\mathrm{sc}}\Psi^{-\infty,0},\text{ and }(\widehat{K_{\tau_{0}}})_{\operatorname{ff}}(\tau)=\chi_{\tau_{0}}(\tau)\Pi_{\tau_{0}},$
@@
This follows directly from $(\widehat{P_{V}+K_{\tau_{0}}})_{\operatorname{ff}}(\tau_{0})=(\widehat{P_{V}})_{\operatorname{ff}}(\tau_{0})+\Pi_{\tau_{0}}$, since $\Pi_{\tau_{0}}$ is exactly projection onto the kernel of the self-adjoint operator $(\widehat{P_{V}})_{\operatorname{ff}}(\tau_{0})$. ∎
@@
With the lemma, we can now conclude the proof
@@
The first two terms may both be bounded by $G^{\prime\prime}P_{V}u$ for $G^{\prime\prime}$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(G)\cup{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(G^{\prime})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime\prime})$ by choosing ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(K_{\tau_{0}})$ sufficiently small. Moreover, $P_{V}-P_{V_{+}}=-(V-V_{+})$ and therefore by (8.1) the third term is controlled by $\lVert u\rVert_{-N,\ell-1}$ for any $N$, and we obtain overall that there is $C>0$ such that, if all the terms below are finite, then we have an estimate
@@
Iterating this estimate allows us to drop the $r-1$ term on the right.
@@
  • $\operatorname{fibeq}$ the fiber equator, page 3.14
@@
  • $\hat{N}_{\operatorname{ff}}(A)=\hat{A}_{\operatorname{ff}}$ and $\hat{N}_{\operatorname{ff},r}(A)$ the indicial operators, page 4.6

### v2

@@ 8.2. With bound states
In this section, we prove Theorem 8.3 and Theorem 8.4. We make appropriate adjustments to the above propagator construction to include the possibility that there are bound states of the Hamiltonian $\Delta+m^{2}+V_{\pm}$ with (negative) energy bigger than 00. Such states appear as elements in the kernel of $\hat{N}_{\operatorname{ff}}(P_{V})(\tau)$ for $\tau\in(-m,m)$.
@@
$\hat{N}_{\operatorname{ff},+}(P_{V})(\tau)=\mbox{ indicial operator of $P_{V}$ at $\mathrm{NP}$},$
while $\hat{N}_{\operatorname{ff},-}(P_{V})(\tau)$ is the indicial operator at $\mathrm{SP}$. Similarly we write
@@
$(\hat{N}_{\operatorname{ff},\pm}(P_{V})(\tau)w=0\implies w\in\mathcal{S}(\mathbb{R}^{n}),$
and thus by self-adjointness of $\hat{N}_{\operatorname{ff},\pm}(P_{V})(\tau)$,
(8.15) $\tau\in\operatorname{Ell}_{\operatorname{ff},\pm}(P_{V})\iff\ker(\hat{N}_{\operatorname{ff},\pm}(P_{V})(\tau))=\{0\}.$
@@
It is possible that there are finitely many points $\tau_{0}\in(-m,m)$ such that $\hat{N}_{\operatorname{ff},\pm}(P_{V})(\tau_{0})$ is not invertible, and the remainder of this section proves estimates near such $\tau_{0}\in W^{\perp}$.
@@
$E_{\pm}(\lambda(\tau_{0}))\coloneqq\ker\left(\hat{N}_{\operatorname{ff},\pm}(P_{V})(\tau_{0})\colon H_{\mathrm{sc}}^{s,\ell}(\operatorname{ff})\longrightarrow H_{\mathrm{sc}}^{s-2,\ell}(\operatorname{ff})\right).$
@@
Assume that $V$ is asymptotically static of order $0,1$ (8.18) holds. Then for any $\epsilon>0$, there exists $Q\in{}^{3\mathrm{sc}}\Psi^{0,0}(\mathbb{R}^{n+1})$ such that
@@
Let $\tau_{0}\in\mathcal{B}_{V,+}$, i.e. let $\tau_{0}\in(-m,m)$, $\tau_{0}\neq 0$ and $E_{+}(\lambda(\tau_{0}))\neq\{0\}$. For any $Q\in{}^{3\mathrm{sc}}\Psi^{0,0}(\mathbb{R}^{n+1})$ with $\tau_{0}\subset\operatorname{Ell}_{\operatorname{ff}}(Q)$ and ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)$ sufficiently small, for any, $G\in{}^{3\mathrm{sc}}\Psi^{0,0}(\mathbb{R}^{n+1})$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$, then for $s,\ell,N,M\in\mathbb{R}$, the estimates in Proposition 8.7 hold.
@@
(8.24) $K_{\tau_{0}}\in{}^{3\mathrm{sc}}\Psi^{-\infty,0},\text{ and }\hat{N}_{\operatorname{ff}}(K_{\tau_{0}})(\tau)=\chi_{\tau_{0}}(\tau)\Pi_{\tau_{0}},$
@@
This follows directly from $\hat{N}_{\operatorname{ff}}(P_{V}+K_{\tau_{0}})(\tau_{0})=\hat{N}_{\operatorname{ff}}(P_{V})(\tau_{0})+\Pi_{\tau_{0}}$, since $\Pi_{\tau_{0}}$ is exactly projection onto the kernel of the self-adjoint operator $\hat{N}_{\operatorname{ff}}(P_{V})(\tau_{0})$. ∎
@@
With the lemma, we can now conclude the proof.
@@
The first two terms may both be bounded by $G^{\prime\prime}P_{V}u$ for $G^{\prime\prime}$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(G)\cup{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(G^{\prime})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime\prime})$ by choosing ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(K_{\tau_{0}})$ sufficiently small. Moreover, $P_{V}-P_{V_{+}}=-(V-V_{+})$ and since $V$ is asymptotically static of order $0,1$, the third term is controlled by $\lVert u\rVert_{-N,\ell-1}$ for any $N$, and we obtain overall that there is $C>0$ such that, if all the terms below are finite, then we have an estimate
@@
Iterating this estimate allows us to drop the $\ell-1$ term on the right.
@@
  • ${}^{3\mathrm{sc}}\mathcal{R}$ the $3\mathrm{sc}$-radial set, page 3.14
  • $\operatorname{fibeq}$ the fiber equator, page 3.15
@@
  • $\hat{N}_{\operatorname{ff}}(A)=\hat{A}_{\operatorname{ff}}$ and $\hat{N}_{\operatorname{ff},\ell}(A)$ the indicial operators, page 4.6